OpenStudy (anonymous):

n which situation is the maximum possible work done? a.when angle between force and displacement is 0 b.when angle between force and displacement is 180 c.when angle between force and displacement is 45 d.when angle between force and displacement is 90

OpenStudy (souvik):

\(W=F*d*cos\theta\) where F=force ,d=distance and \(\theta\) is the angle between F and d so the W would be maximum when \(cos\theta\)=1 of \(\theta\)=0